Solve the system of linear equations:
6x+4y=18
9x+6y=27
If the system has infinitely many solutions, express your answer in the form x=x and y as a function of x
x=
y=

Answer by Elomeht(22)

Both equations are actually saying the same thing:
1. Divide the first equation by 2: 3x + 2y = 9
2. Divide the second equation by 3: 3x + 2y = 9
So the most we can say is:
x = x; y = (9 - 3x)/2